OpenStudy (anonymous):

Bethany can mow her family’s lawn in 4 hours. Her brother Colin can mow the lawn in 3 hours. Which equation can be used to find x, the amount of time it would take for Bethany and Colin to mow the lawn together?

OpenStudy (***[isuru]***):

hi, \[\frac{ 1 }{ time \ by \ bethany } + \frac{ 1 }{ time \ by \ colin } = \frac{ 1 }{ time together }\] \[\frac{ 1 }{4 } + \frac{ 1 }{ 3 } = \frac{ 1 }{ time together }\] \[\frac{ 7 }{12 } = \frac{ 1 }{time \ together }\] \[time \ together = \frac{ 12 }{ 7 }\] \[time together = 1 . 71 \ hours\\] hope this will help ya!!

OpenStudy (anonymous):

thank you @***[ISURU]***